OpenStudy (anonymous):

A bakery want to make a dessert out of chocolate and almonds. If the cost of chocolate is $.75 a pound and the cost of almonds is $.50 a pound, then how many pounds of chocolate should be used to make 30 pounds of the dessert that costs the bakery $.65 a pound?

OpenStudy (zzr0ck3r):

a+b=30 75a+50b=30*75